OpenStudy (anonymous):

From the sum of 5 – 9i and -4 + 7i, subtract 14 – 3i

OpenStudy (azureilai):

Add 5-9i and -4+7i together first, so 5-9i-4+7i Then combine like terms and subtract 14-3i

OpenStudy (anonymous):

so its -13-5i right

OpenStudy (azureilai):

1-2i-(14-3i) -13+i You forgot to distribute the negative one when subtracting. because -2i-(-3i)=i

OpenStudy (anonymous):

thanks can you help me with this majorly hard problem If (x + yi) + (2(x + yi) + 3i) = 9, what is x + yi?

OpenStudy (azureilai):

Just consider x+yi as another number. Problems like these try to use different symbols to confuse you. The equation comes out to x+yi+2x+2yi+3-=9 Now you can combine like terms.

OpenStudy (anonymous):

3yi+3x+3=9?

OpenStudy (azureilai):

Yes but put the 3 in front first. Standard form is a+bi so rewrite it to 3x+3yi+3=9 Im not sure if your teacher wants you to even simplify it down further, but you can also say 3x+3yi-6=0

OpenStudy (anonymous):

she want it simplified a little further

OpenStudy (azureilai):

You can also rewrite it to 3(x+yi-2) = 0

OpenStudy (anonymous):

think you can help with one more problem